a+3-1*3+2-4+0*1=0,a=_?
Answer by colliefan(242)

a+3-1*3+2-4+0*1=0,a=_?
This requires you to know the order of operations. If you don't have any parentheses, the operations should be done basically in the order of most to least complicated and then from left to right.
Exponents are done first, then multiplication and division, then addition and subtraction.
I'll do the problem, putting parentheses around each operation that needs to be done next. There are no exponents, so mult and division come first.
a+3-1*3+2-4+0*1=0
a+3-(1*3)+2-4+0*1=0
a+3-(3)+2-4+0*1=0
a+3-3+2-4+0*1=0
a+3-3+2-4+(0*1)=0
a+3-3+2-4+(0)=0
a+3-3+2-4=0
a+(3-3)+2-4=0
a+(0)+2-4=0
a+2-4=0
a+(2-4)=0
a+(-2)=0
a-2=0
Now, what you do to one side, you must do to the other:
a-2=0
a-2+2=0+2
a+0=2
a=2